The trip of the Amazon River in canoes and back to the mouth of the Amazon took 12 hours. Going upstream, the canoes traveled at 5km/h, but downstream they traveled 10 km/h faster. how far upstream did they go?
Answer by mananth(16946)

let distance be x miles
upstream speeed = 5km/h
down stream speed= 15km/h
..
time upstream + time downstream = total time
..
speed = distance/speed
..
x/5 + x/15 = 12
3x+x /15 =12
4x/15=12
*15
4x=15*12
4x=180
x= 45 miles
m.ananth@hotmail.ca